# MVP-Shot: Multi-Velocity Progressive-Alignment Framework for Few-Shot Action Recognition

### v1

@@ III-C Progressive Semantic-Tailored Interaction
For some challenging action categories with diverse speeds, only utilizing single-velocity temporal features may lead to less satisfactory video classification results. Thus, we propose a novel Progressive Semantic-Tailor Interaction (PSTI) module to capture temporal features from different velocity scales, where each temporal feature captures unique perspectives of temporal representations. In FSAR, with limited support videos, the learned velocity features tend to deviate from the underlying motion semantic without semantic information constraints. We argue that textual and visual information compensates for each other, and using linguistic prior knowledge promotes the visual features to pay more attention to semantic-related features. To this end, PSTI consists of multiple Semantic-Tailor Interaction blocks. Each block progressively utilizes the output text information of the last stage to enhance current semantic-related velocity features, and then each stage outputs corresponding velocity action features for velocity-aware matching.
@@
As shown in Figure 3, the Semantic-Tailor Interaction block (STI) consists of two sub-networks: Temporal Relation Network and Channel Correction Network. These two networks compensate each other to inject semantic information into visual features, which allows the interaction between semantic information and visual features on the temporal and channel domains, respectively. Take an STI block as an example. In the STI block that processes information at the velocity scale $n$, the input visual features and textual features are $\bm{X}_{\mathrm{in}}^{n}\in{\mathbb{R}}^{T\times C}$ and $\bm{q}_{\mathrm{in}}^{n}\in{\mathbb{R}}^{1\times C}$, respectively.
Temporal Relation Network:
Inspired by the prompt methods in NLP [62], we concatenate the visual features with the corresponding text feature, and feed them together into the Temporal Relation Network to obtain the enhanced features. Concretely, given the text feature $\bm{q}_{\mathrm{in}}^{n}\in{\mathbb{R}}^{1\times C}$ and the visual features $\bm{X}_{\mathrm{in}}^{n}\in{\mathbb{R}}^{T\times C}$ at the velocity scale $n$, we obtain new sequence features $\hat{\bm{X}^{n}}$ by concatenating the text features and the visual features.
Then we feed the extended sequence features $\hat{\bm{X}^{n}}$ into the remaining transformer layers, which consist of layer normalization (LN), Multi-Headed Self-Attention (MSA), and MLP block to allow the interaction between the semantic token and the visual tokens along the temporal dimension. Concretely, the interim interaction output $\dot{\bm{X}^{n}}$ is computed as
@@
where $\mathrm{MSA}^{h}$ is a single attention head $h$ in MSA, $\bm{F}_{\mathrm{q}}^{h}=\mathrm{LN}(\hat{\bm{X}^{n}})\bm{W}_{\mathrm{q}}^{h}$ is the query, $\bm{F}_{\mathrm{k}}^{h}=\mathrm{LN}(\hat{\bm{X}^{n}})\bm{W}_{\mathrm{k}}^{h}$ is the key, and $\bm{F}_{\mathrm{v}}^{h}=\mathrm{LN}(\hat{\bm{X}^{n}})\bm{W}_{\mathrm{v}}^{h}$ is the value. We concatenate all the attention heads and perform linear projection to obtain the final output $\dot{X^{n}}$. Then, we further apply a feed-forward network to obtain enhanced feature $\bar{\bm{X}^{n}}$ as shown in Fig. 3, given by:
@@
where FFN is composed of two linear projections that are separated by a GELU non-linearity. The feature dimension stays unchanged.
Channel Correction Network:
In addition to taking full account of temporal interaction between the visual features and the text semantic, we believe that taking advantage of text semantic information to perform channel interaction is conducive to multi-velocity feature construction. Thus, we propose a Channel Correction Network that utilizes the channel interaction to correct feature channels and enhance semantic-related velocity features. Specifically, given enhanced visual features $\bar{\bm{X}^{n}}=[\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{T}]$, we first obtain a global visual context vector $\bar{\bm{x}_{\mathrm{g}}^{n}}$ by averaging all visual tokens:
@@
where $\mathrm{MLP}(\cdot)$ represents the MLP layer consisting of two fully connected layers and a ReLU layer. Finally, we add the modulated global vector to all visual tokens so that it can tune the visual features at each channel. The final modulated visual features $\bm{X}_{\mathrm{out}}^{n}\in{\mathbb{R}}^{T\times C}$ can be written as:
@@
In addition, in order to preserve high-level scene information, we design a progressive connection operation to connect the visual outputs of the STI block at different velocity scales and pool features from the raw video input ${\bm{X}_{\mathrm{in}}^{n}}$. Given the interaction feature output $\bm{X}_{\mathrm{out}}^{n}\in{\mathbb{R}}^{T\times C}$, the progressive connection uses temporal fusion to provide $\bm{X}_{\mathrm{in}}^{n+1}\in{\mathbb{R}}^{\frac{T}{2}\times C}$ for the next block, as follows,
@@
Similarly, we adopt the same operation on the query set videos to explore the temporal relationship. However, since the real class information of the query video is unknown during testing, we only input the visual query feature into the Semantic-Tailor Interaction block. For the whole PSTI module, we repeat the above process to recurrently inject velocity-tailored text information and progressively interact with video-text information at different velocity levels.
@@ III-D Multi-Velocity Feature Alignment
where $\mathrm{D}(\bar{\bm{X}^{n}_{\mathrm{s}}},\bar{\bm{X}^{n}_{\mathrm{q}}})$ is the distance between the support visual feature $\bar{\bm{X}^{n}_{\mathrm{s}}}$ and the query visual feature $\bar{\bm{X}^{n}_{\mathrm{q}}}$. $\mathcal{M}$ is the OTAM metric [11] by default in our MVP-Shot. Note that the proposed MVP-Shot is a plug-and-play framework and in the later experimental section, we will insert MVP-Shot into other existing metrics or methods and empirically demonstrate its pluggability. Furthermore, we express the total support-query distance between support and query videos as the weighted sum of the distances obtained by multiple velocity scales:
@@
where ${\alpha}_{n}$ is the coefficient. By default, the maximum velocity scale $N$ is set to $3$ in this paper for considering the balance between effectiveness and efficiency.
To tackle the issue that different action instances may be performed at different motion dynamics, we propose PSTI and MVFA to progressively capture and align multi-velocity action features. By this means, our proposed MVP-shot combines two modules to complement each other to predict query categories more accurately under the few-shot setting.
Finally, we can use the output total support-query distance $\mathrm{D}_{\mathrm{s},\mathrm{q}}$ as logits for classification. Following previous works, we can use a cross-entropy loss $\mathcal{L}_{CE}$ to optimize the model parameters. For few-shot inference, we can leverage the distance values of support-query video pairs in (11)) to classify query samples based on the nearest neighbor rule [39].
TABLE I: Comparison with recent state-of-the-art few-shot action recognition methods (§IV-C) on Kinetics [21], UCF101 [20] and HMDB51 [19]. The experiment settings are conducted under the $5$-way $K$-shot. ”INet-RN50” denotes ResNet-50 pre-trained on ImageNet“-” means the result is not available in published works. The best results are highlighted.
@@
TABLE II: Comparison with recent state-of-the-art few-shot action recognition methods (§IV-C) on SSv2-small [22] dataset. The experiment settings are conducted under the $5$-way $K$-shot. ”INet-RN50” denotes ResNet-50 pre-trained on ImageNet“-” means the result is not available in published works. The best results are highlighted.
@@ IV Experiments
In this section, we first describe four few-shot action recognition datasets ( i.e., Kinetics [21], UCF101 [20], HMDB51 [19], and SSv2-small [22]) and the evaluation protocol (§IV-A). Then we introduce the implementation Details of our MVP-shot (§IV-B). After that, we compare our methods with previous state-of-the-art few-shot action recognition methods on four gold-standard datasets (§IV-C). Finally, we perform several ablation experiments to demonstrate the effectiveness of the proposed MVP-shot (§IV-D).
@@ IV-A Experimental Setup
Dataset. We evaluate the effectiveness of our proposed method on four benchmarks, i.e., Kinetics [21], SSv2-small [22], UCF101 [20], and HMDB51 [19], using $5$-way $1$/$3$/$5$-shot few-shot recognition tasks for all models considered. For Kinetics and SSv2-small, we follow the split setting as in CMN [6], which both randomly select $64$ classes from the original dataset as the train set, $12$ classes as the val set and $24$ classes as the test set. For HMDB51[19], we use the splits from [17], and the dataset has 31 action classes, 10 action classes, and 10 action classes for train/val/test, respectively. The UCF101[20] dataset contains 101 action classes, and we adopt the splits from [17], which splits the dataset into 70/10/21 classes for train/val/test, respectively.
Evaluation. Following the official evaluation protocols [9, 11], we use $5$-way $1$-shot and $5$-way $5$-shot accuracy for evaluation, and report average results over 10,000 tasks randomly selected from test.
@@ IV-B Implementation Details
Network Architecture. For a fair comparison with previous works [11, 14] that use Resnet-50 [65] pre-trained on ImageNet [66] as the backbone, we provide a version of utilizing pre-trained CLIP ResNet-50 as our basic visual encoder. To retain the original pre-trained prior knowledge in the text encoder and reduce the optimization burden, we freeze the text encoder and prevent it from being updated during training.
Network Training. Following previous methods [67], we sparsely and uniformly sample $T$ = $8$ frames to obtain the representation of the whole video. Each frame is scaled to a height of $256$. Each video frame is first resized to 256 × 256, and then these resized video frames are randomly cropped into the size of 224 × 224. In the training phase, we adopt basic data augmentation, such as random horizontal flipping, cropping, and color jitter. Moreover, our framework uses the Adam optimizer with the multi-step scheduler to train our model. The total number of training steps is set to $10$. In the inference pahse, we report the mean video classification accuracy in 10000 episodes randomly sampled from the test sets. For many shot classifications, we follow the prototype paradigm  [39] and calculate the mean features of support videos in each class as the prototypes to classify query videos.
Reproducibility. Our model is implemented in PyTorch and trained on two TITAN RTX GPUs with a $24$GB memory per-card. Testing is conducted on the same machine. To guarantee reproducibility, full code will be released.
@@ IV-C Comparison with SOTA methods
Results on Kintices [21], HMDB51 [19] and UCF101 [20]. Table  I demonstrates our method outperforms other state-of-the-art methods on Kinetics, UCF101, and HMDB51 datasets under the 1-shot to 5-shot setting. Table I demonstrates our method outperforms other state-of-the-art methods on Kinetics, UCF101, and HMDB51 datasets under the 1-shot to 5-shot setting. Specifically, our model significantly improves accuracy in any task setting compared to previous excellent methods (such as TRX [14], HyRSM [7], and MoLo [64]). Compared to CLIP-FSAR  [60], which uses the same CLIP pre-training and temporal alignment metric, our MVP-shot achieves better results in multiple datasets and task settings. Notably, compared to CLIP-FSAR [60], our method brings $3.3$%, $2.4$% performance improvements in the 1-shot task of HMDB51 and Kinetics, and $2.2$%, $1.3$% gains in the 5-shot task of HMDB51 and Kinetics, respectively.
Results on SSv2-small [22]. We also report our MVP-shot model results on SSV2-small datasets on Table II. Our CLIP-RN50 model achieves the best performance in the $3$-shot and $5$-shot tasks compared to all the methods using ResNet-50 as the visual encoder. Compared to CLIP-FSAR [60], our method also has a significant performance improvement. Concretely, our MVP-shot brings $1.0$% accuracy improvements in the 3-shot task, and $1.2$% accuracy gains in the 5-shot task, respectively.
@@
Excellent performance on these datasets reveals that our MVP-shot has strong robustness and generalization for different datasets. In addition, from the experimental results, we can make the following observations. (1) Compared with CLIP-FSAR [60] using the same CLIP pre-training and temporal alignment metric, our MVP-shot achieves better results in multiple datasets and task settings. Specifically, MVP-shot brings $3.3$% and $2.4$% performance improvements in the 1-shot tasks of HMDB51 and Kinetics, respectively. It indicates our MVP-shot further boosts performance by exploring multi-velocity feature matching. (2) We observe that MVP-shot outperforms current state-of-the-art methods (such as TRX [14], HyRSM [7], and MoLo [64]), especially under the $1$-shot and $3$-shot settings. We attribute this to the fact that introducing text semantic cues would be more effective for enhancing semantic-related velocity features when visual information is insufficient.
@@ IV-D Ablation Study
We conduct ablation experiments to validate the capability of the proposed MVP-shot and analyze the contribution of each component. Unless otherwise specified, we adopt the CLIP-RN50 model as the default visual encoder.
Analysis of Each Module in MVP-shot. To validate the contributions of each module in our method, we conduct experiments under $5$-way $1$-shot and $5$-shot settings on the Kinetics [21] datasets. Table III summarizes the effects of each module in MVP-shot. By performing multi-velocity feature alignment (MVFA) (c.f. §III-C), we obtain $2.1$% and $2.5$% $1$-shot and $5$-shot improvements, respectively. This consistent promotion indicates that MVFA can capture pair-wise semantic relevance at different velocity scales to facilitate accurate matching between videos. In addition, the proposed Progressive Semantic-Tailored Interaction (PSTI) module (c.f. §III-D) improves on $1$-shot and $5$-shot by $8.5$% and $3.0$%, respectively, which indicates recurrently injecting velocity-tailored text information significantly boosts performance. We observe that the performance gain is particularly significant in the 1-shot scenario, suggesting that it would be more effective to supplement textual semantic information when visual information is very limited. Moreover, stacking MVFA and PSTI modules can further improve performance, indicating the complementarity between components.
TABLE V: Comparison of different feature interaction methods (§IV-D). We reported the $1$-shot and $5$-shot results on Kinetics [21] dataset. The best results are highlighted.
@@
Analysis of Multi-Velocity Feature Alignment. To demonstrate that multi-velocity feature alignment coordinates with each other to facilitate accurate matching between videos, we conduct the experiments on the $5$-way $1$-shot and $5$-shot tasks of Kinetics [21] by various combinations of multi-velocity feature alignment. The obtained results are reported in Table IV. When the action feature is provided at a single velocity, the corresponding performance is reduced on the $1$-shot and $5$-shot settings. The performance of multi-velocity feature alignment improves as the velocity scales increase. Further, our MVP-shot, which combines all velocity feature alignments, achieves the best performance. Specifically, on the $1$-shot setting, MVP-shot outperforms all single-velocity feature alignments by $1.7$%, $4.0$%, and $8.4$%, respectively. These results demonstrate that our proposed multi-velocity feature alignment can surpass corresponding methods with single-velocity feature alignment, and capture pair-wise semantic relevance at different velocity scales to facilitate coherent matching between videos.
TABLE VI: Generalization performance of MVP-shot with different temporal alignment metrics (§IV-D) on Kinetics [21] in the $5$-way $1$-shot and $5$-way $5$-shot tasks.
@@
Figure 4: $N$-way $1$-shot results of our MVP-shot on Kinetics [21] dataset, CLIP-FSAR and our baseline methods with $N$ varying from $5$ to $10$.
Comparison with Different Feature Interaction Methods. To demonstrate the effectiveness of our proposed Semantic-Tailored Interaction (STI) module and compare the efficacy of various feature interaction methods, we conduct experiments on the 5-way 1-shot and 5-way 5-shot task of Kinetics [21]. Specifically, we replace our STI module with several other feature interaction methods. The transformer[68] consists of the multi-head self-attention and a feed-forward network. The first baseline single-modal transformer indicates the support features and query features doing self-attention on the temporal dimension without text features. Compared with the first baseline, the second multi-modal transformer ( CLIP-FSAR [60] ) stacks the text features to the corresponding support features along the temporal dimension before self-attention. We set all the layers of the transformer to be one. For a fair comparison, the other experimental settings are the same as the proposed MVP-shot. The experimental results are shown in Table  V. In comparison, the proposed STI module outperforms the unimodal transformer by $8.5$% ($90.0$% vs. $81.5$%) and $1.2$% ($93.2$% vs. $92.0$%) on the $1$-shot and $5$-shot settings, respectively. The performance gains indicate that there is complementarity between visual and textual modalities, and semantic information can help visual features to be more discriminative. Compared with the multimodal transformer, the proposed STI module increases $1.7$% ($88.3$% vs. $90.0$%) and $0.9$% ($92.3$% vs. $93.2$%) on the $1$-shot and $5$-shot settings, respectively. Experimental results demonstrate demonstrates that STI module enhances the utilization of textual information more efficiently to capture the semantic-related velocity features. Thus, our proposed module can achieve more robust and more accurate class prototype estimation.
TABLE VII: Influence of different backbones. Quantitative comparison results (§IV-D) on Kinetics [21], UCF101 [20], HMDB51 [19], and SSv2-small [22]. “-” means the result is not available in published works. The best results are highlighted.
@@
Generalization on Different Alignment Metrics. We conduct the experiments using different temporal alignment metrics on the $5$-way $1$-shot and $5$-way $5$-shot tasks of Kinetics [21] to demonstrate our model generalizability. We adopt different temporal alignment metrics, including OTAM [11] and Bi-MHM [7]. As shown in Table VI, our method can adapt to any temporal alignment metric, and achieve great performance. Moreover, irrespective of the temporal alignment metric employed, our framework still achieves impressive performance gains compared to the naive baselines, indicating the superiority of our method.

### v4

@@ III-C Progressive Semantic-Tailored Interaction
For some challenging action categories with diverse speeds, only utilizing single-velocity temporal features may lead to less satisfactory video classification results. Thus, we design a novel Progressive Semantic-Tailor Interaction (PSTI) module to capture temporal features from different velocity scales, where each temporal feature captures unique perspectives of temporal representations. In FSAR, with limited support videos, the learned velocity features tend to deviate from the underlying motion semantic without semantic information constraints. We argue that textual and visual information compensates for each other, and using linguistic prior knowledge promotes the visual features to focus more on semantic-related features. To this end, PSTI consists of multiple Semantic-Tailor Interaction blocks. Each block progressively utilizes the output text information of the last stage to enhance current semantic-related velocity features, and then each stage outputs corresponding velocity action features for velocity-aware matching.
@@
As shown in Figure 3, the Semantic-Tailor Interaction block (STI) comprises two sub-networks: Temporal Relation Network and Channel Correction Network. These two networks compensate each other by injecting semantic information provided by class names into visual features, which perform feature interaction on the temporal and the channel dimensions, respectively. Take an STI block as an example. In the STI block that processes information at the velocity scale $n$, the input visual features and textual features are $\bm{X}_{\mathrm{in}}^{n}\in{\mathbb{R}}^{T\times C}$ and $\bm{q}_{\mathrm{in}}^{n}\in{\mathbb{R}}^{1\times C}$, respectively.
(a) Temporal Relation Network: Inspired by the prompt methods in NLP [62], we concatenate the visual features with the corresponding text feature, and feed them together into the Temporal Relation Network to obtain the enhanced features. Concretely, given the text feature $\bm{q}_{\mathrm{in}}^{n}\in{\mathbb{R}}^{1\times C}$ and the visual features $\bm{X}_{\mathrm{in}}^{n}\in{\mathbb{R}}^{T\times C}$ at the velocity scale $n$, we obtain new sequence features $\hat{\bm{X}^{n}}$ by concatenating the text features and the visual features.
Then we feed the extended sequence features $\hat{\bm{X}^{n}}$ into the remaining transformer layers, which consist of Multi-Headed Self-Attention (MSA), and MLP layer to allow the interaction between the semantic token and the visual tokens along the temporal dimension. Concretely, we compute the interim interaction output $\dot{\bm{X}^{n}}$ as
@@
where $\mathrm{MSA}^{h}$ is a single attention head $h$ in MSA, $\bm{F}_{\mathrm{q}}^{h}=\mathrm{LN}(\hat{\bm{X}^{n}})\bm{W}_{\mathrm{q}}^{h}$ is the query, $\bm{F}_{\mathrm{k}}^{h}=\mathrm{LN}(\hat{\bm{X}^{n}})\bm{W}_{\mathrm{k}}^{h}$ is the key, and $\bm{F}_{\mathrm{v}}^{h}=\mathrm{LN}(\hat{\bm{X}^{n}})\bm{W}_{\mathrm{v}}^{h}$ is the value. We concatenate all the attention heads and perform linear projection to obtain the final output $\dot{X^{n}}$. Then, as shown in Fig. 3, we further apply FFN to obtain enhanced feature $\bar{\bm{X}^{n}}$, given by:
@@
where FFN consists of two linear projections separated by a GELU activation function.
(b) Channel Correction Network: In addition to fully accounting for temporal interaction between the visual features and the text semantic, we believe that taking advantage of text semantic information to perform channel interaction is conducive to multi-velocity feature construction. Thus, we propose a Channel Correction Network that utilizes the channel interaction to correct feature channels and enhance semantic-related velocity features. Specifically, given enhanced visual features $\bar{\bm{X}^{n}}=[\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{T}]$, a global visual context vector $\bar{\bm{x}_{\mathrm{g}}^{n}}$ is firstly obtained by averaging all visual tokens:
@@
where $\mathrm{MLP}(\cdot)$ denotes the MLP layer, which comprises two fully connected layers and a ReLU activation layer. Finally, we add the modulated global vector to all visual tokens so that visual features pay more attention to semantic information. The final modulated visual features $\bm{X}_{\mathrm{out}}^{n}\in{\mathbb{R}}^{T\times C}$ can be written as:
@@
In addition, we design progressive connection operation to connect the visual outputs of the STI block at different velocity scales and pool features from the raw video input ${\bm{X}_{\mathrm{in}}^{n}}$, so as to preserve high-level scene information. Given the interaction feature output $\bm{X}_{\mathrm{out}}^{n}\in{\mathbb{R}}^{T\times C}$, the progressive connection uses temporal fusion to provide $\bm{X}_{\mathrm{in}}^{n+1}\in{\mathbb{R}}^{\frac{T}{2}\times C}$ for the next block, as follows,
@@
Similarly, to explore the temporal relationships, we adopt the same operation on the query set videos but do not utilize category names. For the whole PSTI module, we repeat the above process to recurrently inject velocity-tailored text information and progressively interact with video-text information at different velocity levels.
@@ III-D Multi-Velocity Feature Alignment
where $\mathrm{D}(\bar{\bm{X}^{n}_{\mathrm{s}}},\bar{\bm{X}^{n}_{\mathrm{q}}})$ is the distance between the query visual feature $\bar{\bm{X}^{n}_{\mathrm{q}}}$ and the support visual feature $\bar{\bm{X}^{n}_{\mathrm{s}}}$. $\mathcal{M}$ denotes the OTAM temporal metric [11] by default in our MVP-Shot. The proposed MVP-Shot is designed as a plug-and-play framework. We will incorporate MVP-Shot into other existing metrics in the following experimental section, and empirically demonstrate its pluggability. Furthermore, we express the total support-query distance between support and query videos as the weighted sum of the distances calculated across multiple velocity scales:
@@
where ${\alpha}_{n}$ is the coefficient. The maximum velocity scale $N$ is set to $3$ by default in this paper for considering the balance between effectiveness and efficiency.
To tackle the issue that different action instances may be performed at different motion dynamics, we propose PSTI and MVFA to progressively capture and align multi-velocity action features. By this means, our proposed MVP-shot combines two modules to complement each other to make more accurate query predictions under the few-shot setting.
Finally, we can use the output total query-support distance $\mathrm{D}_{\mathrm{s},\mathrm{q}}$ as logits for further classification. Following previous works [7, 9, 63], we can optimize the model parameters using a cross-entropy loss $\mathcal{L}_{CE}$. For few-shot inference, we can obtain final prediction results based on the nearest neighbor rule [39] by utilizing the distance values of support-query video pairs in Eq.(11).
TABLE I: Comparison results with recent state-of-the-art few-shot action recognition methods (§IV-C) on Kinetics [21], UCF101 [20] and HMDB51 [19]. We conduct experiments under the $5$-way $K$-shot settings. We highlight the best results. ”INet-RN50” denotes ResNet-50 visual encoder.
@@
TABLE II: Comparison results with recent state-of-the-art few-shot action recognition methods (§IV-C) on SSv2-small [22] dataset. We conduct experiments under the $5$-way $K$-shot settings. We highlight the best results. ”INet-RN50” denotes ResNet-50 visual encoder.
@@ IV Experiments
In this section, we first introduce four few-shot action recognition datasets ( i.e., UCF101 [20], Kinetics [21], HMDB51 [19], and SSv2-small [22]) and the evaluation protocol (§IV-A). Then we describe the implementation details of our MVP-shot (§IV-B). Following that, we benchmark our methods against previous state-of-the-art few-shot action recognition approaches on four widely used datasets (§IV-C). Finally, to demonstrate the effectiveness of the proposed MVP-shot, we conduct several ablation experiments(§IV-D).
@@ IV-A Experimental Setup
Dataset. We evaluate the effectiveness of our proposed MVP-shot on four benchmarks, i.e., Kinetics [21], SSv2-small [22], UCF101 [20], and HMDB51 [19]. We evaluate all models on different datasets using $5$-way $1$/$3$/$5$-shot few-shot recognition tasks. For SSv2-small and Kinetics, we utilize the split setting as in CMN [6], which both randomly select $64$ classes from the original dataset as the train set, $12$ classes as the val set and $24$ classes as the test set. The UCF101[20] dataset consists of 101 action classes, and we utilize the splits as in [17], where there are 70, 10, and 21 classes for train/val/test, respectively. For HMDB51[19], we utilize the splits from [17], and the dataset has 31 action classes, 10 action classes, and 10 action classes for train/val/test, respectively.
Evaluation. Following the official evaluation protocols [9, 11], we use $5$-way $1$-shot, $3$-shot and $5$-shot accuracy for evaluation. We randomly select $10000$ episodic tasks from test and report average results.
@@ IV-B Implementation Details
Network Architecture. We utilize pre-trained CLIP ResNet-50 as our basic visual encoder to ensure a fair comparison with previous methods [11, 14], which choose Resnet-50 [65] pre-trained on ImageNet [66] as the visual encoder. In order to make full use of the pre-trained prior knowledge in CLIP and alleviate the optimization burden, we freeze the text encoder and prevent it from being updated during training.
Network Training. Following previous methods [67], we sparsely and uniformly sample $T$ = $8$ frames of the raw video sequences to encode the frame-level video representation. Each video frame is scaled to a height of $256$. Then we randomly crop these resized video frames into the size of 224 × 224. In the training phase, following previous methods [7, 60], we employ basic data augmentation, such as color jitter, random horizontal flipping, and cropping. Moreover, we use the Adam optimizer with the multi-step scheduler to train our model. By default, we set the total number of training steps to $10$. In the inference phase, we use the average of the 10000 episodes as the result. For many shot classifications, we follow the prototype paradigm [39] and acquire the mean features of support videos within each class to serve as prototypes, so as to calculate the distance between prototypes and query videos for classification. To guarantee reproducibility, full code will be released.
@@ IV-C Comparison with SOTA methods
Results on Kintices [21], HMDB51 [19] and UCF101 [20]. AS shown in Table  I, we compare our MVP-shot with current state-of-the-art few-shot action recognition on three standard datasets. Table  I demonstrates our method outperforms other state-of-the-art methods on Kinetics, HMDB51, and UCF101 datasets under different shot settings. Specifically, our model has achieved significant improvements in accuracy under any task setting compared to previous excellent methods (such as MoLo [63], HyRSM [7], and GgHM [9]). Compared to CLIP-FSAR [60], which makes use of semantic information to perform cross-modal interaction, our MVP-shot achieves better results in multiple datasets in that our method performs multi-velocity feature matching. Notably, our method achieves an increase of $3.3$%, $2.4$% performance improvements on HMDB51 and Kinetics datasets under the 1-shot task, and achieves an increase of $2.2$%, $1.3$% performance improvements on HMDB51 and Kinetics datasets under the 5-shot task compared to CLIP-FSAR [60], respectively.
Results on SSv2-small [22]. We also report our MVP-shot model results on SSV2-small datasets on Table II. Our MVP-shot achieves better performance in the $3$-shot and $5$-shot tasks than all other state-of-the-art methods on SSV2-small datasets. Compared to CLIP-FSAR [60], our method also has a significant performance improvement. Concretely, our MVP-shot brings $1.0$% performance improvements in the 3-shot task, and $1.2$% performance improvements under the 5-shot setting, respectively.
@@
Our MVP-shot achieves the best performance on all datasets, which reveals that our MVP-shot has strong generalization across diverse datasets. In addition, we can observe the following from the experimental results: (1) Compared with CLIP-FSAR [60], which makes use of semantic information to perform frame-level feature matching, our MVP-shot achieves better performance in multiple task settings and datasets. Specifically, MVP-shot brings $3.3$% and $2.4$% performance improvements on HMDB51 and Kinetics under the 1-shot setting, respectively. It indicates our MVP-shot further boosts performance by exploring multi-velocity feature matching. (2) We observe that MVP-shot achieves better performance than current state-of-the-art methods (such as HyRSM [7], TRX [14], and MoLo [63]), especially under the $1$-shot and $3$-shot settings. This indicates that incorporating textual semantic information would be more effective for enhancing semantic-related velocity features when visual information is insufficient.
@@ IV-D Ablation Study
We conduct ablation experiments to substantiate the efficacy of our proposed MVP-shot and dissect the contribution of each component. Unless explicitly stated otherwise, we adopt the CLIP Resnet-50 model as our visual encoder by default.
Analysis of Each Module in MVP-shot. To validate the impacts of each module in our MVP-shot, we conduct experiments on the Kinetics [21] under $5$-way $1$-shot and $5$-shot settings. Table III summarizes the effects of each module in MVP-shot. By performing multi-velocity feature alignment (MVFA) (c.f. §III-C), we obtain $2.1$% and $2.5$% performance gains under $1$-shot and $5$-shot settings, respectively. The consistent promotion indicates that MVFA can capture pair-wise semantic relevance at different velocity scales to facilitate accurate matching between videos. In addition, the proposed Progressive Semantic-Tailored Interaction (PSTI) module (c.f. §III-D) improves by $8.5$% and $3.0$% on $1$-shot and $5$-shot settings, respectively, which indicates recurrently injecting velocity-tailored text information significantly boosts performance. We observe a notable performance improvement, particularly in the 1-shot scenario, indicating that the supplementation of textual semantic information proves more effective when visual information is severely limited. Moreover, stacking MVFA and PSTI modules can compensate for each other to further improve performance.
TABLE V: Comparison of different feature interaction methods (§IV-D). We reported the $1$-shot and $5$-shot results on Kinetics [21] dataset.
@@
Analysis of Multi-Velocity Feature Alignment. To demonstrate that multi-velocity feature alignment coordinates with each other to facilitate accurate matching between videos, we conduct the experiments on Kinetics [21] under the $5$-way $1$-shot and $5$-shot settings by various combinations of multi-velocity feature alignment. We report the obtained results in Table IV. When the action feature is provided at a single velocity, the corresponding performance is reduced on the $1$-shot and $5$-shot settings. The performance of multi-velocity feature alignment improves as the velocity scales increase. Further, our MVP-shot, which combines all velocity feature alignments, achieves the best performance. Specifically, MVP-shot outperforms all single-velocity feature alignments by $1.7$%, $4.0$%, and $8.4$% on the $1$-shot setting, respectively. These results demonstrate that our proposed multi-velocity feature alignment can surpass corresponding methods with single-velocity feature alignment, and capture pair-wise semantic relevance at different velocity scales to facilitate coherent matching between videos.
Figure 4: $N$-way $1$-shot results of our MVP-shot on Kinetics [21] dataset, our baseline methods and CLIP-FSAR with $N$ varying from $5$ to $10$.
Comparison with Different Feature Interaction Methods. To evaluate the efficacy of different feature interaction methods and demonstrate the superiority of our proposed Semantic-Tailored Interaction (STI) module, we perform comparative analysis on Kinetics [21] under the 5-way 1-shot and 5-way 5-shot settings. Specifically, we replace our STI module with several other feature interaction methods. The transformer[68] consists of the multi-head self-attention multi-head self-attention mechanisms and a feed-forward network. The single-modal transformer indicates the support features and query features performing self-attention on the temporal dimension without text features. Compared with the first baseline, the second multi-modal transformer ( CLIP-FSAR [60] ) stacks the support features to corresponding textual features along the temporal dimension before self-attention. All the layers of the transformer are set to be one. All other experimental settings remain consistent with those of the proposed MVP-shot to ensure a fair comparison. We report the experimental results in Table  V. In comparison, the proposed STI module outperforms the unimodal transformer by $8.5$% ($90.0$% vs. $81.5$%) and $1.2$% ($93.2$% vs. $92.0$%) under the $1$-shot and $5$-shot settings on Kinetics dataset, respectively. The observed performance gains suggest complementarity between textual and visual modalities, wherein semantic information augments the discriminative capacity of visual features. Compared with the multimodal transformer, the proposed STI module increases $1.7$% ($88.3$% vs. $90.0$%) and $0.9$% ($92.3$% vs. $93.2$%) under the $1$-shot and $5$-shot settings on Kinetics dataset, respectively. Experimental results demonstrate that STI module enhances the utilization of textual information more efficiently to capture the semantic-related velocity features. Thus, our proposed module can achieve more robust and more accurate class prototype estimation.
TABLE VI: Generalization performance of MVP-shot with different temporal alignment metrics (§IV-D) on Kinetics [21] under the $5$-way $1$-shot and $5$-way $5$-shot settings.
@@
TABLE VII: Influence of different backbones. Quantitative comparison results (§IV-D) on Kinetics [21], UCF101 [20], HMDB51 [19], and SSv2-small [22]. We highlight the best results.
@@
${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ We utilize different temporal alignment metrics to conduct experiments on Kinetics [21] under the $5$-way $1$-shot and $5$-shot settings to demonstrate our model ${}_{\!}$ Different temporal alignment metrics are adopted, including OTAM [11] and Bi-MHM [7]. As shown in Table VI, our MVP-shot is adaptable to any temporal alignment metric and achieves great performance. Moreover, irrespective of employed temporal alignment metric, compared to the naive baselines, our framework still achieves significant performance gains, indicating the superiority of our MVP-shot.